OpenStudy (anonymous):

Differentiate A=s ^{2}, A=1/2 bh, A=1/2hb _{1}+1/2hb _{2}, A=pi r ^{2}, A=2 pi rh+ 2pi r^{2}

OpenStudy (anonymous):

\[Differentiate A=s ^{2}, A=1/2 bh, A=1/2hb _{1}+1/2hb _{2}, A=\pi r ^{2}, A=2 \pi rh+ 2\pi r^{2}\]